OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given point and having the given slope. C(0, 4), m = 0

OpenStudy (anonymous):

if the gradient is zero its a horizontal line that is parallel to the x axis so there wouldn't be a constant or an x or gradient shown for that matter therefore y=4

OpenStudy (anonymous):

Eq of line in slope point form is given as: \[y-y_1= m(x-x_1)\] \[y-4= 0(x-0)\rightarrow y-4= 0 \rightarrow y=4 \] Hence y=4 is the standard form the equation of the line. @caitlin_lynette_munoz

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

what about Indicate in standard form the equation of the line passing through the given points. E(-2, 2), F(5, 1)

OpenStudy (anonymous):

well, then you'd have to find the gradient first, which is rise/run or y1-y2/x1-x2 so 2-1/(-2)-5=-1/7 then substitute in one of the points into the formula equation: y=mx+c y=-1/7x+c 1=-1/7(5) +c c=12/7 you've got your constant, and your gradient, now fill in the equation: _=_ _ +/- _

OpenStudy (anonymous):

gradient means slope

OpenStudy (anonymous):

so would it be y=(-1/7)x+(12/7)

OpenStudy (anonymous):

yes